OpenStudy (anonymous):

-absolute value of x+2 (end of absolute value) +5 greater than or equal to 3

OpenStudy (diyadiya):

\[-|x+2|+5 \ge 3\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

r u there??

OpenStudy (anonymous):

subtract 5 from both sides

OpenStudy (anonymous):

i did that but that but i still dont understand!

OpenStudy (roadjester):

what are you supposed to do? You just give us a problem, but no instructions.

OpenStudy (anonymous):

i dont know! i dont understand what to do!

OpenStudy (roadjester):

what are your insturctions? where did the problem come from?

OpenStudy (roadjester):

are you solving for x, are you graphing? what?

OpenStudy (anonymous):

my math book! it just says solve

OpenStudy (roadjester):

ok, SOLVE FOR WHAT!!! what is the section title?

OpenStudy (anonymous):

section title is:absolute value equations and inequalities

OpenStudy (roadjester):

Ok, I think it wants the range of solutions. Do you have a final answer? If you do, this will be much easier, just don't give it to me until I'm done.

OpenStudy (anonymous):

no its just multiple choice

OpenStudy (roadjester):

Well, that's ok I guess. Here I go.

OpenStudy (anonymous):

multiple choice answers are: A.\[3le x \le5\]. B.\[-1le x \le1\]. C.\[-4\le x \le0\] or D.\[-2\le x \le5\]

OpenStudy (anonymous):

A.\[3\le x \le\]

OpenStudy (roadjester):

\[-\left| x+2 \right|+5\ge3\] subtract 5 from both sides \[-\left| x+2 \right|\ge-2\] divide by negative \[\left| x+2 \right|\le2\] \[-2\le x+2 \le 2\] subtract 2 from all three sides \[-4 \le x \le 0\]

OpenStudy (anonymous):

thank u!

OpenStudy (roadjester):

Do I get a medal? lol

OpenStudy (anonymous):

sure! lol!

OpenStudy (roadjester):

YAY!

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

\[x^-2\] divided by 13xy^-2